OpenStudy (anonymous):

Can Someone Explain this to me? 5x-3y-16=0 4x-5y-5=0 Solution Set: x| 2 |5 5(2)-3-16 5(5)-3y-16=0 y|-2 |3 10-3y-16 25-36-16=0 3y=16-10 36=16-25 3y=6 3y=9 3 3 y=-2 3 3 y=3 x|5|0 4(5)-5y-5=0 -5y-5=0 y|3|-1 20-5y=0 -5y=5 -5

OpenStudy (espex):

While it appears to be a system of equations, what are you asked to do and what are you stuck on?

OpenStudy (anonymous):

everything i dont know how they get the absolute value of 5x-3y-16=0 which is x| 2 |5 y|-2 |3

OpenStudy (espex):

I do not believe that is an absolute value, it appears to be a separation for the values.

OpenStudy (anonymous):

x| | y| | this is blank and you have to find the absolute value of this 5x-3y-16=0 to get it. can you help me understand this?

OpenStudy (anonymous):

dd u try what i said??....if u put x=2 in that eqn...what value of y do u get???

OpenStudy (espex):

You do not have to find the absolute value, if that is what they wanted they would have given you something like |x|=? You are confusing yourself by worrying about the ABS, don't. As @Sourav16 suggested, just use a plug and chug method.

OpenStudy (anonymous):

yea..its a trial and error prob...

OpenStudy (anonymous):

can you help me solve this 1.) 2x-3y=2 4x+5y=26?

OpenStudy (espex):

How would you like to solve it, substitution or elimination?

OpenStudy (anonymous):

like the example that have given.

OpenStudy (espex):

The example you are given is just illustrating how you can take a table of values for x, plug them in, and calculate the value of y. Do you have a list of x and/or y values?

OpenStudy (anonymous):

i have no list of x and y values, just that. i have to find x and y also

OpenStudy (espex):

Well that will be easy then. Take your first equation, 2x-3y=2, pick three (3) values for 'x' and plug them in one at a time, solve for y.

OpenStudy (anonymous):

first select a variable...if u select x...keep y unchanged in the equation...then put values of x...starting from 1....ultimately...u'll get a value of x which satisfies the equation...let me solve one of the eqns... \[2x-3y=2 => 2x-2=3y =>y =2x-3/3\] now...put x=1...u'll get y=0 put x=2...u'll get y=2/3....but thats fractional....u cant select this value of y... put x= 4, u get y= 2..... remember..if u want to do it easily...put the values of x in that last step of the equation i simplified....dat'll be a lot easier and quicker....

OpenStudy (espex):

Your example equation should read \[y =\frac{2}{3}x-\frac{2}{3}\]

OpenStudy (anonymous):

can you guys give me the x here then ill do it from there x+y-6=0 3x-5y-34=0

OpenStudy (anonymous):

getting x is now the only thing i dont know.

OpenStudy (espex):

You make it up. Pick some values, plug in these and fill in the resulting y. x | y ------ -2 | -1 | 0 | 1 | 2 |

OpenStudy (anonymous):

2(4)-3y=2 8-3y=2 -3y=2-8 -3y=-6 x|4| -3 -3 = y|2| i need another one but is this right?

OpenStudy (espex):

Yes, that is done perfectly. Excellent work.

OpenStudy (anonymous):

another is 2(7)-3y=2 14-3y=2 -36=2-14 -3y=-12 -3 -3 = y=4

OpenStudy (espex):

Again, correct.

OpenStudy (anonymous):

what if y has no value like this? x+y-6=0 3x-5y-34=0

OpenStudy (espex):

A variable that stands alone has an invisible 1, so 'y' is the same as '1y' if it is easier for you to use.

OpenStudy (anonymous):

4x+5y=26 4(4)+5y=26 4(9)+5y=26 5y=26-16 5y=26-36 x|4|9 5y=10 y=2 5y=10 y=2 y|2|2| am i correct? and the signs also

OpenStudy (espex):

On your 9/2 values you dropped a - 26-36= -10

OpenStudy (anonymous):

x|4|9 so its y|2|-2|

OpenStudy (espex):

Precisely.

OpenStudy (anonymous):

can you help me find the x and y of this? x+y-6=0 i cant seem to get it

OpenStudy (espex):

It is the same process, you pick a value for x and solve for y, or pick a value for y and solve for x.